OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? Question 3 options: (1, 7) (5, 0) (2, -1) (0, 5)

OpenStudy (anonymous):

i keep trying but I cant figure it out!

OpenStudy (anonymous):

I get -6x=-15 +5

OpenStudy (anonymous):

same concept as before, substitute each individual point (1,7) into both equations to see if it works.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

tell me ur answer afterwards, I will confirm it for u.

OpenStudy (anonymous):

the first one isn't right... right.. i got -21 +5

OpenStudy (anonymous):

no wait I lost!

OpenStudy (anonymous):

give me a second.. working on it

OpenStudy (anonymous):

kk.

OpenStudy (anonymous):

Example using the point (1,7) and substituting into the first equation. y=-3x +5 7=-3(1) + 5 7=-3 +5 7=2 That is clearly wrong, try the rest yourself.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

option two isn't correct because 0 doesnt = 5

OpenStudy (anonymous):

yep, option two is wrong.

OpenStudy (anonymous):

option 3 is wrong because -1 doesn't equal 9

OpenStudy (anonymous):

try option 3 again.

OpenStudy (anonymous):

u may have made a mistake plusing/minusing.

OpenStudy (anonymous):

wait i confused ... option 4 worked for me.. For option three I got -1= 9

OpenStudy (anonymous):

(2,-1) y = -3x + 5 and y = 2x + 5? -1= -3(2) + 5 -1 = 2(2) + 5 -1 = -6 + 5 -1 = 4 + 5 -1 = -1 -1 = 9 yes, u are right.

OpenStudy (anonymous):

Option D is the answer.

OpenStudy (anonymous):

ok YEAH! THANKS SO MUCH. .... I sorry... I am so bad at these... thanks so much! YOU get a medal!

OpenStudy (anonymous):

haha it's ok, as long as u learnt how to do it :)

OpenStudy (anonymous):

THanks!